# Med-R2: Crafting Trustworthy LLM Physicians via Retrieval and Reasoning of Evidence-Based Medicine

### v4

@@ 3.3 Evidence Applying
3.4 Effect Assessment
Our assessment of evidence is centered around two key aspects: the accuracy of the chain-of-thought (CoT) generator in responding to queries and the stability of evidence document retrieval throughout different stages. Algorithm 2 delineates the evidence evaluation process and the iterative retrieval loop for the query corpus within the pipeline, where we determine the number of iterations for optimization termination based on our assessment of these two factors. This dual focus ensures that when evaluating the LLM physician (target model) on medical tasks, the retrieved evidence documents provided are both effective and robust.
4 Experiments and Results
In this section, we first describe the details of our experimental setup (Section 4.1). We then outline the baseline methods we use for comparison (Section 4.2) and our experimental results to prove the effectiveness of Med-R2 (Section 4.3).
Table 2: Comparison of Med-R2 with baselines. The best and second best are in bold and underlined.
@@ 4 Experiments and Results
GPT-4o 84.95 79.28 74.60 77.20 84.44 80.09
Claude3.5-Sonnet 83.20 72.54 68.80 76.40 83.90 76.97
DeepSeek-V3 80.93 77.82 74.30 73.60 87.67 78.86
@@
PMC-LLaMA-7B 38.65 35.30 32.40 52.26 26.87 37.10
MEDITRON-7B 46.76 42.78 36.83 58.32 43.60 45.66
PMC-LLaMA-13B 40.38 33.07 37.55 54.63 46.85 42.50
MEDITRON-70B 55.05 52.21 53.50 72.84 71.45 61.01
@@
Qwen2.5-7B Direct Response 22.58 39.14 28.77 54.16 44.45 37.82
Vanilla RAG 58.28 64.78 32.78 55.68 54.97 53.30
Fine-Tuning 78.56 84.94 47.43 54.40 56.09 64.28
LLM-AMT 70.44 72.39 43.86 56.02 54.43 59.43
Med-R2 80.67 81.23 49.18 56.74 70.96 67.76
@@
Fine-Tuning 76.53 85.24 47.38 42.78 42.92 58.97
LLM-AMT 52.91 66.63 45.08 44.60 45.39 50.92
Med-R2 77.01 84.16 52.33 52.75 45.19 62.29
LLaMA2-13B Direct Response 25.84 23.00 29.68 40.71 43.57 32.56
Vanilla RAG 36.97 22.15 35.14 52.67 45.81 38.55
Fine-Tuning 39.51 26.65 43.35 55.37 45.70 42.12
LLM-AMT 38.65 25.03 39.87 52.36 48.67 40.92
Med-R2 38.06 29.34 41.96 57.12 48.89 43.07
@@
Fine-Tuning 55.87 85.52 47.64 54.29 80.41 64.75
LLM-AMT 51.48 77.11 43.09 62.42 78.78 62.58
Med-R2 54.34 80.19 48.36 68.32 84.03 67.05
@@
Fine-Tuning 25.57 89.97 66.17 66.69 82.40 66.16
LLM-AMT 19.38 88.08 68.33 68.61 82.06 65.29
Med-R2 24.43 90.01 69.09 68.36 84.95 67.37
@@
Fine-Tuning 87.17 86.21 71.65 74.72 79.46 79.84
LLM-AMT 79.18 68.59 70.12 79.05 80.74 75.54
Med-R2 86.37 84.58 73.36 78.24 82.82 81.07
4.1 Experimental Setup
Model Details. We employ the latest open-sourced LLMs from LLaMA [44, 9] and Qwen [51] series as our target language models for evaluation. We assessed models including Qwen2.5-7B, LLaMA3.1-8B, LLaMA2-13B, Qwen2.5-14B, Qwen2.5-32B and LLaMA3.1-70B, scaling from 7B to 70B. The default setting of context window for our main experiments is 4K, with an in-depth scaling analysis presented in Section 5.
Datasets. We have selected five medical datasets including PubMedQA [22], MedQA-USMLE, MedQA-MCMLE [21], MedMCQA [35], and MMLU-Med [16], and use accuracy as evaluation metrics. More details about the datasets can be found in Section C.1.
Implementation. We constructed the medical knowledge corpus by establishing the FAISS vector library [24]. Experiments related to model training were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 3 epochs. We conducted all training and evaluation experiments on NVIDIA RTX H800 GPUs with 80G memory.
4.2 Baselines
We compare Med-R2 with the following baselines: (1) Frontier Models contain GPT-4o [17], Claude3.5-Sonnet [3] and DeepSeek-V3 [30]. (2) Open-Sourced Medical Models include PMC-LLaMA-7B, PMC-LLaMA-13B [47], MEDITRON-7B and MEDITRON-70B [7]. (3) The simplest baseline is Direct Response, where the model answer medical questions directly without the aid of external knowledge bases or dataset fine-tuning. (4) Vanilla RAG [28] utilizes raw queries for evidence searching, and the retrieved documents are then directly integrated into the generation process without any further manipulation. (5) Fine-Tuning leverages medical datasets to further train the model under supervised conditions. Here we have employed two distinct strategies: within-dataset fine-tuning, where the datasets for training and evaluation are derived from different parts of the same data corpus, and cross-dataset fine-tuning, where the model is fine-tuned on one medical dataset (e.g., MedMCQA) and then evaluated on different datasets (e.g., PubMedQA). (6) LLM-AMT [46] is a dedicated process tailored for biomedical question answering, which includes typical modules such as query augmenter, hybrid retriever, knowledge refiner, etc. Details are discussed in Section C.2.
4.3 Main Results
We conduct evaluations to validate the efficiency of Med-R2 across different open-sourced models. We summarize the observations below.
Med-R2 is effective across different models. Table 2 shows that the incorporation of external knowledge bases significantly enhances the model’s ability to address medical queries, where even the most basic vanilla RAG method depicts an average enhancement of 17.92% over direct responses. Furthermore, Med-R2 provides an added layer of the enhancement by adhering to the EBM process, which outperforms all baselines across various benchmarks, achieving an average improvement of 35.10% over the direct response strategy. Notably, for lightweight models such as Qwen-2.5-7B and LLaMA3.1-8B, Med-R2 demonstrates increases of 79.16% and 77.06% respectively. We surmise that this is due to the fact that while lightweight models inherently lack comprehensive domain-specific medical knowledge, they possess the capability to efficiently read and identify information from external medical documents. Consequently, effective augmentation from external knowledge bases substantially bolsters the models’ capacity to tackle medical-domain questions. Compared with medical-specific models of the same parameter scale, models equipped with our Med-R2 demonstrate superior performance in medical tasks. Moreover, LLaMA3.1-70B + Med-R2 has the potential to surpass frontier models under medical scenarios, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80%. In order to evaluate Med-R2 in real clinical settings, we have performed experiments on several datasets related to real-world clinical scenarios as well, detailed in Section C.4. Additionally, we have also ablate the components of Med-R2  in Table 10 to further identify the contribution of each module.
Med-R2 shows superiority compared to fine-tuning. From Table 2, we find that Med-R2 stands out as the only approach among those leveraging external knowledge bases that surpasses the average performance of fine-tuning methods. Specifically, Med-R2 exhibits nearly equivalent performance to fine-tuning strategies in within-dataset training, yet it significantly outperforms in cross-dataset training, achieving an enhancement of 8.59% and an overall capability improvement of 3.32%. One contributing factor is that during within-dataset fine-tuning, the training and testing datasets are of the same origin, thus a model trained on homogeneous data should logically achieve substantial performance gains on the test set. Conversely, in cross-dataset fine-tuning, the heterogeneity between the training and testing datasets more rigorously assesses the model’s ability to generalize within the same domain. Under these circumstances, the utilization of a comprehensive external medical knowledge base and the effective retrieval and extraction of pertinent information becomes particularly crucial. This phenomenon is even more pronounced in Table 8 when evaluating models in real clinical scenarios where the complexity and variability of patient cases require a higher degree of information retrieval and evidence synthesis. Additionally, considering that fine-tuning necessitates additional training time and computational resources, Med-R2 emerges as a more efficient approach for enhancing the model’s performance on medical domain-related issues.
5 Ablations and Analysis Across Scales
Previously in Section 4, we have established the efficacy of Med-R2 in improving models’ medical performances. Here we further analyze the impact of model scale and context window length on Med-R2 , and provide rationale for incorporating the CoT sequences of models into the retrieval process starting from the second iteration.
Integrating CoT at the onset may bring adverse effects. We ablate the components of query reformulator $q_{EBM}\left[+q_{CoT}\right]$ through integrating the chain-of-thought (CoT) sequences generated by models into the initial evidence retrieval phase, which is denoted as Med-R2-CoT. Figure 4 presents a grouped bar chart illustrating the medical task performance of Med-R2-CoT and all baselines included in our main experiments. It is observed that Med-R2-CoT exhibits a decline in performance compared to Med-R2, with the disparity increasing as the model parameter scale decreases. Notably, at the 7-8 billion parameter level (e.g., Qwen2.5-7B and LLaMA3.1-8B), the performance of Med-R2-CoT is even inferior to that of the vanilla RAG strategy. We hypothesize that one contributing factor is that the lightweight models’ less solid grasp of medical knowledge. As a result, in the absence of supplementary external medical knowledge bases, models of these scales may struggle to align the direction of their thinking with the original query, and thus, the generated CoT sequences may even negatively impact the retrieval effectiveness. However, this phenomenon is somewhat mitigated as the model parameter scale increases.
Figure 4: Grouped bar chart depicting the average evaluation results of baselines from Table 2 and Med-R2-CoT. Compared to Med-R2, Med-R2-CoT involves the immediate incorporation of models’ CoT sequence into the query reformulator for retrieval during the initial round.
@@ 5 Ablations and Analysis Across Scales
Effect of context window scale on model’s performance. We compared the performance of Med-R2 across models with varying parameter scales and different context window lengths on medical tasks, as shown in Table 9, and then plotted heatmaps illustrating the percentage improvement of Med-R2 over direct responses, as depicted in Figure 5. It reveals that Med-R2 exhibits an optimal context window length for enhancing the model’s medical performance, which increases with the growth of model parameter size. Concurrently, the enhancement of Med-R2 follows a trend of initial decline followed by an increase with the escalation of model scale. Specifically:
@@
Figure 5: The percentage increase of performance achieved by Med-R2 over direct responses across various context window sizes and model scales.
We hypothesize that as the context length increases, the role of Med-R2’s reranker diminishes since most retrieved evidence documents are fed into the same context window of the model. At this point, lightweight models, particularly those around 8B, which may not have a solid grasp of medical knowledge, and overly long evidence sequences could reduce the model’s efficiency in extracting key information, potentially leading to the generation of hallucinations and adversely affecting the model’s medical performance. However, the addition of other modules, such as the query reformulator, can help improve the precision of knowledge retrieval, thereby mitigating this negative impact to some extent. In contrast, models with 32B and above scales exhibit greater robustness, and the impact of increasing window length on Med-R2’s effectiveness is less pronounced.
6 Discussion
In this study, we follow the Evidence-Based Medicine (EBM) process to design a novel LLM physician framework, which effectively leverages the retrieval, filtering, and reasoning processes inherent to EBM. Experiments demonstrate that Med-R2 holds superior advantages over existing strategies, while also reducing the substantial computational costs associated with model training. Furthermore, our analysis of model scale and context window size also highlights the scaling capabilities of Med-R2. By scaling the model parameters and adjusting the context window length, we demonstrate the robustness of Med-R2 in achieving optimal performance across a broad spectrum of medical tasks. However, there are some limitations in our work: (1) Due to the professional nature and rapid evolution of medical knowledge, the medical knowledge corpus we construct may not ensure complete coverage of medical information, potentially limiting the upper bound of retrieval precision. (2) During the fine-grained reranking phase, to balance the computational cost and effectiveness, we employ a lightweight proxy model to calculate the usefulness score of the current evidence document, yet it does not fully represent the performance tendencies of the target model during actual evaluating.

### v5

@@ 3.3. Evidence Applying
3.4. Effect Assessment
Our assessment of evidence is centered on the following aspect: the stability of evidence document retrieval across different stages. Algorithm 2 delineates the evidence evaluation process and the iterative retrieval loop for the query corpus within the pipeline, where we determine the number of iterations for optimization termination based on our assessment of such factor. This ensures that when evaluating the LLM physician (target model) on medical tasks, the retrieved documents provided are effective and robust.
Params: Coarsely reranked documents $\mathcal{S}^{c}$, finely reranked documents $\mathcal{S}^{f}$
Define $E(d)$: embedding of document $d$
1:  $\mathcal{S}_{\mathcal{Q}-top}^{f},\mathcal{Q}^{CoT}\leftarrow[\,],\,[\,]$
2:  for each query $q_{i}$ in $\mathcal{Q}$ do
3:   for $t=1,2,\ldots,T$ do
5:    /* Step 1: CoT Generation */
6:    Select top $k$ documents from $\mathcal{S}_{i}^{f^{(t)}}$ referring to $w$:
$\mathcal{S}_{i-top}^{f^{(t)}}\leftarrow\text{SelectTopK}(\mathcal{S}_{i}^{f^{(t)}},k,w)$
7:    Generate chain-of-thought sequence based on $q_{i}$ and $\mathcal{S}_{i-top}^{f^{(t)}}$:
$q^{CoT^{(t)}}_{i}\leftarrow M_{CoT}(q_{i},\mathcal{S}_{i-top}^{f^{(t)}})$
8:    /* Step 2: Evidence Assessment */
9:    Compute semantic stability of docs in consecutive iterations:
[…]
10:    /* Step 3: Termination Condition Evaluation */
11:    if $\vec{\mu}^{(t)}_{i}<\delta$ then
12:     $\mathcal{S}_{\mathcal{Q}-top}^{f}.append(\mathcal{S}_{i-top}^{f^{(t)}})$, $\mathcal{Q}^{CoT}.append(q^{CoT^{(t)}}_{i})$
13:     Break
14:    end if
15:   end for
16:   $\mathcal{S}_{\mathcal{Q}-top}^{f}.append(\mathcal{S}_{i-top}^{f^{(T)}})$, $\mathcal{Q}^{CoT}.append(q^{CoT^{(T)}}_{i})$
17:  end for
18:  Return Selected documents $\mathcal{S}_{\mathcal{Q}-top}^{f}$ and CoT sequences $\mathcal{Q}^{CoT}$
4. Experiments and Results
Table 2. Comparison of Med-R2 with baselines. The best and second best are in bold and underlined.
Model Method MedQA-USMLE MedQA-MCMLE MedMCQA PubMedQA MMLU-Med NEJMQA MedXpertQA RareArena-RDC RareArena-RDS Avg.
@@ 4. Experiments and Results
GPT-4o 84.95 79.28 74.60 77.20 84.44 72.20 30.37 73.76 47.50 69.31
Claude3.5-Sonnet 83.20 72.54 68.80 76.40 83.90 68.52 21.31 68.46 50.78 65.99
DeepSeek-V3 80.93 77.82 74.30 73.60 87.67 74.23 24.16 75.09 45.72 68.73
@@
PMC-LLaMA-7B 38.65 35.30 32.40 52.26 26.87 35.63 11.02 28.96 21.67 31.42
MEDITRON-7B 46.76 42.78 36.83 58.32 43.60 38.80 11.74 35.68 26.50 37.89
PMC-LLaMA-13B 40.38 33.07 37.55 54.63 46.85 40.60 11.44 39.86 28.98 37.04
MEDITRON-70B 55.05 52.21 53.50 72.84 71.45 66.97 18.05 74.81 46.58 56.83
@@
LLaMA3.1-8B Direct Response 31.16 41.45 30.02 36.17 37.12 50.41 14.90 41.85 22.16 33.91
Vanilla RAG 55.80 59.38 35.91 47.10 43.54 52.01 15.50 56.23 35.06 44.51
Fine-Tuning 76.53 85.24 47.38 42.78 42.92 52.48 15.54 55.38 38.37 50.73
LLM-AMT 52.91 66.63 45.08 44.60 45.39 53.07 16.22 60.20 40.29 47.16
Med-R2 77.01 84.16 52.33 52.75 45.19 55.69 17.79 62.90 44.82 54.74
Qwen2.5-14B Direct Response 50.01 65.23 42.85 56.93 71.60 45.63 11.06 43.17 26.58 45.89
Vanilla RAG 54.88 75.60 42.06 60.38 79.46 45.69 12.28 54.00 28.05 50.27
Fine-Tuning 55.87 85.52 47.64 54.29 80.41 46.67 12.30 66.32 43.66 54.75
LLM-AMT 51.48 77.11 43.09 62.42 78.78 47.44 13.73 60.03 38.81 52.54
Med-R2 54.34 80.19 48.36 68.32 84.03 48.03 14.45 72.80 46.30 57.43
Qwen2.5-32B Direct Response 16.23 87.07 66.44 68.66 80.19 43.59 15.06 59.67 31.33 52.03
Vanilla RAG 19.33 89.30 67.63 67.06 83.85 46.63 17.49 66.10 40.59 55.33
Fine-Tuning 25.57 89.97 66.17 66.69 82.40 52.58 15.71 70.20 45.08 57.15
LLM-AMT 19.38 88.08 68.33 68.61 82.06 46.82 17.90 66.75 45.40 55.93
Med-R2 24.43 90.01 69.09 68.36 84.95 50.40 18.54 72.51 55.36 59.29
LLaMA3.1-70B Direct Response 46.43 58.36 62.33 66.81 71.33 65.43 23.43 72.59 43.54 56.69
Vanilla RAG 62.66 77.91 66.63 68.78 76.95 66.40 25.95 77.70 44.88 63.10
Fine-Tuning 87.17 86.21 71.65 74.72 79.46 68.11 24.18 76.41 47.23 68.35
LLM-AMT 79.18 68.59 70.12 79.05 80.74 68.90 26.53 77.00 45.87 66.22
Med-R2 86.37 84.58 73.36 78.24 82.82 70.18 26.84 78.60 49.42 70.04
4.1. Experimental Setup
Model Details. We employ the open-sourced LLMs from LLaMA (Touvron et al., 2023; Dubey et al., 2024) and Qwen (Yang et al., 2024) series as our target models for evaluation. We assessed models including LLaMA3.1-8B, Qwen2.5-14B, Qwen2.5-32B and LLaMA3.1-70B, scaling from 8B to 70B. The default setting of context window for our main experiments is 4K, with an in-depth scaling analysis presented in Section 5.
Datasets. We have selected eight medical datasets including PubMedQA (Jin et al., 2019), MedQA-USMLE, MedQA-MCMLE (Jin et al., 2020), MedMCQA (Pal et al., 2022), MMLU-Med (Hendrycks et al., 2021), NEJMQA (Katz et al., 2024), MedXpertQA (Zuo et al., 2025), and RareArena (THUMedInfo, 2025), covering both standard and real-world clinical scenarios. We use accuracy as evaluation metrics. More details about the datasets can be found in Section C.1.
Implementation. We constructed the medical knowledge corpus by establishing FAISS vector library (Johnson et al., 2019). Experiments related to model training were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 3 epochs. We conducted all training and evaluation experiments on NVIDIA RTX H800 GPUs with 80G memory.
4.2. Baselines
We compare Med-R2 with the following baselines: (1) Frontier Models contain GPT-4o (Hurst et al., 2024), Claude3.5-Sonnet (Anthropic, 2024) and DeepSeek-V3 (Liu et al., 2024). (2) Open-Sourced Medical Models include PMC-LLaMA-7B, PMC-LLaMA-13B (Wu et al., 2024), MEDITRON-7B and MEDITRON-70B (Chen et al., 2023). (3) The simplest baseline is Direct Response, where the model answer medical questions directly without the aid of external knowledge bases or dataset fine-tuning. (4) Vanilla RAG (Lewis et al., 2020) utilizes raw queries for evidence searching, and the retrieved documents are then directly integrated into the generation process without any further manipulation. (5) Fine-Tuning leverages medical datasets to further train the model under supervised conditions. Here we employ two strategies: within-dataset fine-tuning, where the datasets for training and evaluation are derived from different parts of the same data corpus, and cross-dataset fine-tuning, where the model is fine-tuned on one medical dataset (e.g., MedMCQA) and then evaluated on different datasets (e.g., PubMedQA). (6) LLM-AMT (Wang et al., 2024) is a dedicated process tailored for biomedical question answering, which includes typical modules such as query augmenter, hybrid retriever, knowledge refiner, etc. Details are discussed in Section C.2.
4.3. Main Results
We have performed evaluations to validate the efficiency of our Med-R2 on open-sourced models across different parameter scales. The main results of baselines and Med-R2 are demonstrated in Table 2, and we summarize the observations below.
Med-R2 is effective across different models. Table 2 shows that the incorporation of external knowledge bases significantly enhances the model’s ability to address medical queries, where even the most basic vanilla RAG method depicts an average enhancement of 13.10% over direct responses. Furthermore, Med-R2 provides an added layer of the enhancement by adhering to the EBM process, which outperforms all baselines across benchmarks, achieving an average improvement of 28.10% over the direct response strategy. Notably, for lightweight models such as LLaMA3.1-8B, Med-R2 demonstrates increases of 61.43%. We surmise that this is due to the fact that while lightweight models inherently lack comprehensive domain-specific medical knowledge, they possess the capability to efficiently read and identify information from external medical documents. Consequently, effective augmentation from external knowledge substantially bolsters the models’ capacity to tackle medical-domain questions. Compared with medical-specific models of the same parameter scale, models equipped with Med-R2 demonstrate superior performance in medical tasks. Moreover, LLaMA3.1-70B + Med-R2 has the potential to surpass frontier models under medical scenarios, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.05%, 6.14% and 1.91%.
Med-R2 shows superiority compared to fine-tuning. From Table 2, we find that Med-R2 stands out as the only approach among those leveraging external knowledge that surpasses the average performance of fine-tuning methods. Specifically, Med-R2 exhibits nearly equivalent performance to fine-tuning strategies in within-dataset training, yet it significantly outperforms in cross-dataset training, achieving an enhancement of 7.61% and an overall capability improvement of 4.55%. One contributing factor is that during within-dataset fine-tuning, the training and testing datasets are of the same origin, thus a model trained on homogeneous data could achieve substantial performance gains on the test set. Conversely, in cross-dataset fine-tuning, the heterogeneity between the training and testing datasets more rigorously assesses the model’s ability of generalization. Under these circumstances, the utilization of a comprehensive external medical knowledge base and the effective retrieval and extraction of pertinent information becomes particularly crucial. Additionally, considering that fine-tuning necessitates additional training time and computational resources, Med-R2 emerges as a more efficient approach for enhancing the model’s performance on medical domain-related issues.
5. Ablations and Analysis Across Scales
We further analyze the impact of model scale and context window length on Med-R2 , and provide rationale for incorporating the CoT sequences of models into the retrieval process starting from the second iteration. Additionally, we have also ablate the components of Med-R2 in Table 3 to identify the contribution of each module.
Integrating CoT at the onset may bring adverse effects. We ablate the components of query reformulator $q_{EBM}\left[+q_{CoT}\right]$ through integrating the chain-of-thought (CoT) sequences generated by models into the initial evidence retrieval phase, which is denoted as Med-R2-CoT. Figure 6 presents the medical task performance of Med-R2-CoT and all baselines included in our main experiments. It is observed that Med-R2-CoT exhibits a decline in performance compared to Med-R2, with the disparity increasing as the model parameter scale decreases. Notably, at the 8 billion parameter level (e.g., LLaMA3.1-8B), the performance of Med-R2-CoT is even inferior to that of the vanilla RAG strategy. We hypothesize that one contributing factor is that the lightweight models’ less solid grasp of medical knowledge. As a result, in the absence of external medical knowledge, models of these scales may struggle to align the direction of their thinking with the original query, and thus, the generated CoT sequences may even negatively impact the retrieval effectiveness. However, this phenomenon is somewhat mitigated as the model parameter scale increases.
Figure 6. Average evaluation results of baselines from Table 2 and Med-R2-CoT. Compared to Med-R2, Med-R2-CoT involves the immediate incorporation of models’ CoT sequence into the query reformulator for retrieval during the initial round.
Figure 7. The percentage increase of performance achieved by Med-R2 over direct responses across various context window sizes and model scales.
Table 3. Module analysis of Med-R2 . We sequentially integrate various modules onto the vanilla RAG systems to conduct comparative analyses. We highlight the optimal performance values for models with single and dual component additions.
@@ 5. Ablations and Analysis Across Scales
+ Query Reformulator 62.47 72.85 41.66 48.72 43.98 53.94
+ Evidence Reranker 68.62 75.74 43.84 49.97 43.76 56.34
+ CoT Generator 59.87 65.63 39.45 47.71 43.61 51.25
Med-R2 (ours) 77.01 84.16 52.33 52.75 45.19 62.29
@@
+ Query Reformulator 54.93 78.96 45.43 64.85 81.73 65.18
+ Evidence Reranker 55.03 78.83 45.97 65.08 81.56 65.29
+ CoT Generator 54.62 76.72 43.86 62.13 80.02 63.47
+ Query Reformulator, CoT Generator 54.77 80.58 47.61 66.98 82.21 66.43
+ Evidence Reranker, CoT Generator 54.81 81.79 47.87 67.48 82.95 66.98
Med-R2 (ours) 54.34 80.19 48.36 68.32 84.03 67.05
@@
+ Query Reformulator 20.14 89.42 68.72 68.41 83.97 66.13
+ Evidence Reranker 20.83 89.88 68.88 68.47 83.99 66.41
+ CoT Generator 19.78 88.69 67.94 67.72 83.92 65.61
+ Query Reformulator, Evidence Reranker 23.04 89.25 70.35 68.89 84.46 67.20
Med-R2 (ours) 24.43 90.01 69.09 68.36 84.95 67.37
@@
+ Query Reformulator 70.82 80.64 69.72 71.82 73.56 73.31
+ Evidence Reranker 73.96 80.78 70.87 74.64 75.28 75.11
+ CoT Generator 68.08 78.65 68.85 70.09 73.72 71.88
Med-R2 (ours) 86.37 84.58 73.36 78.24 82.82 81.07
@@
Effect of context window scale on model’s performance. We compared the performance of Med-R2 across models with varying parameter scales and different context window lengths on medical tasks, and then plotted heatmaps illustrating the percentage improvement of Med-R2 over direct responses, as depicted in Figure 7. It reveals that Med-R2 exhibits an optimal context window length for enhancing the model’s medical performance, which increases with the growth of model parameter size. Concurrently, the enhancement of Med-R2 follows a trend of initial decline followed by an increase with the escalation of model scale. Specifically:
@@
We hypothesize that as the context length increases, the role of Med-R2’s reranker diminishes since most retrieved documents are fed into the same context window of the model. At this point, lightweight models, particularly those around 8B, which may not have a solid grasp of medical knowledge, and overly long sequences could reduce the model’s efficiency in extracting key information, potentially leading to the generation of hallucinations and adversely affecting the model’s medical performance. However, the addition of other modules, such as the query reformulator, can help improve the precision of knowledge retrieval, thereby mitigating this negative impact to some extent. In contrast, models with 32B and higher scales exhibit greater robustness, and the impact of increasing the window length on Med-R2’s effectiveness is less pronounced.
Module Impact Analysis. We have analyzed the contributions of various modules in Med-R2 to the model’s performance in the medical domain using a default context window size of 4K, sequentially incorporating components into the vanilla RAG framework. As shown in Table 3, we decompose the modules into three components: query reformulator, evidence reranker, and CoT generator. Overall, the evidence reranker contributed the most to models’ performances among the individual components. When combined with the query reformulator, the performance gains were even more pronounced, showing a synergistic effect. The addition of the CoT generator further enhanced the model’s ability to effectively utilize retrieved medical documents, providing substantial added value.
6. Discussion
In this study, we follow the Evidence-Based Medicine (EBM) process to design a novel LLM physician framework, which effectively leverages the retrieval, filtering, and reasoning processes inherent to EBM. Experiments demonstrate that Med-R2 holds superior advantages over existing strategies, while also reducing the substantial computational costs associated with model training. Furthermore, our analysis of model scale and context window size also highlights the scaling capabilities of Med-R2. By scaling the model parameters and adjusting the context window length, we demonstrate the robustness of Med-R2 in achieving optimal performance across a broad spectrum of medical tasks.